OpenStudy (daniellelovee):

Divide (x^3+2-5)÷(x-3) Please explain this and how you would solve it

OpenStudy (daniellelovee):

@mathmate

OpenStudy (mathmate):

@Daniellelovee Hey! Is this a long division issue, or factorization? It looks like there will be remainders!

OpenStudy (mathmate):

Is it \(\dfrac{x^3+2x-5}{x-3}\), or \(\dfrac{x^3+2x^2-5}{x-3}\)?

OpenStudy (daniellelovee):

I don't think you can factor it

OpenStudy (daniellelovee):

ok how did you get there

OpenStudy (daniellelovee):

(x^3+2-5)÷(x-3) this is the equation, the 2 does not have an x

OpenStudy (mathmate):

I am not sure about the question. The coefficient 2 does not have a power of x to go with it. I am just checking what the question is like, so far.

OpenStudy (daniellelovee):

thats what Im confused about because if the 2 had an x I would factor it and then divide normally but since it doesn't im just confused

OpenStudy (daniellelovee):

and there is no lest common factor either

OpenStudy (mathmate):

so that would be \(\dfrac{x^3+2-5}{x-3}=\dfrac{x^3-3}{x-3}\) then

OpenStudy (mathmate):

That is why I thought something is missing.

OpenStudy (mathmate):

Even if we put 2x or 2x^2, the numerator does not have rational factors.

OpenStudy (daniellelovee):

so the answer is x once you divide right?

OpenStudy (daniellelovee):

because the x-3 cancels out

OpenStudy (mathmate):

\(\dfrac{x^3+2-5}{x-3}=\dfrac{x^3-3}{x-3}=x^2+3x+9~with~remainder~24\)

OpenStudy (mathmate):

Not sure if that's what you need. Can you double check the question, the +2-3 looks "fishy".

OpenStudy (mathmate):

* +2-5

OpenStudy (mathmate):

Well, if the numerator does not factorize, then the x-3 cannot cancel out. The best we can do is to divide and state the remainder.

OpenStudy (daniellelovee):

lol nvm it was probably a typo since this was a practice question before the quiz

OpenStudy (daniellelovee):

I just wanted to make sure I wasn't the only one that thought that this question was odd.

OpenStudy (mathmate):

I think there must be a type, we don't usually write x^3+2-5! lol